# Video-Foley: Two-Stage Video-To-Sound Generation via Temporal Event Condition For Foley Sound

### v2

@@ IV-D2 Subjective Evaluation
We conducted a human listening test to assess the perceptual quality of the generated audio in relation to the input video. 20 participants were asked to score the audio on a five-point Likert scale based on four criteria: Material / Action / Temporal Alignment, and Audio Quality. Semantic Alignment was divided into two categories to evaluate how well the sound matches the material type and action nuance of the sound events in the video. We provided guidelines and video examples to clearly distinguish Material and Action Alignment from Temporal Alignment during evaluation. The evaluation survey consisted of 12 questions covering different material-action types. Specifically, we excluded ‘None’ from the dataset’s 18 material categories and selected six ‘{material}-scratch’ cases where the sound characteristics significantly change by scratching actions. These cases include plastic-scratch, rock-scratch, dirt-scratch, drywall-scratch, gravel-scratch, and grass-scratch. In addition, we selected six ‘{material}-hit’ cases from the remaining material categories where the sound characteristics notably change by hitting actions. These cases include carpet-hit, ceramic-hit, metal-hit, water-hit, wood-hit, and leaf-hit. To standardize the length of the sample videos and control evaluator fatigue, we trimmed each video to 4 seconds from the starting point. Each question presented the ground truth audio and the audio generated by Video-Foley, SyncFusion, Diff-Foley, CondFoleyGen, and AudioLDM in a random order. Since CondFoleyGen does not support text prompts, audio prompts were used for all models to ensure a fair comparison. The Mean Opinion Score (MOS) and its 95% confidence interval were calculated.
Figure 7: Performance of Video2RMS (left) and RMS2Sound (right) for different numbers of RMS bins. w/o_rms: without RMS condition (Text-to-Audio[6]), cont.: continuous RMS, no discretization.
@@ V-A Analysis on Video2RMS
Table I demonstrates the performance of the Video2RMS model. Our proposed model is compared to random choice (lower bound), the regression approach, and the scores of discretized ground-truth RMS (upper bound due to information loss). In the classification setting, the model significantly outperforms the regression baseline across all metrics, validating our decision to approach RMS prediction as a classification problem. As illustrated in Figure 6, the model effectively predicts the RMS curve for sparse audio events, avoiding the shortcut to predict silence as observed in the regression baseline. The relatively low accuracy in Acc±$\pm$1dB is because it prioritizes predicting a realistic RMS curve over matching the exact magnitude bin. Finally, the label smoothing improves performance, improving both E-L1 and prediction accuracy.
TABLE II: Performance of the proposed Video-Foley and other video-to-sound models on Greatest Hits testset. $av$: audio-video paired prompt used, +: same CLAP model for train and evaluation. Regarding train data, †: VGGSound[33] ($\sim$0.4k hr), $\ddagger$: VGGSound, AudioSet[34] ($\sim$1.3k hr), otherwise: Greatest Hits trainset ($\sim$6 hr).
@@
SpecVQGAN[8] 101.0 579 - 0.0273 - 0.323 6.42 - -
Diff-Foley[10] 87.0 550 2.11($\pm$0.11) 0.0363 1.86($\pm$0.14) 0.403 4.61 1.78($\pm$0.13) 2.38($\pm$0.17)
@@
CondFoleyGenav[17] 42.2 381 3.10($\pm$0.13) 0.0326 1.93($\pm$0.13) 0.572 1.01 2.36($\pm$0.16) 2.79($\pm$0.17)
SyncFusion[12] 65.9 335 3.10($\pm$0.13) 0.0307 3.10($\pm$0.19) +0.631 4.50 3.04($\pm$0.18) 3.22($\pm$0.19)
Video-Foley (Ours) 27.2 187 3.93($\pm$0.12) 0.0126 4.40($\pm$0.11) 0.644 0.80 3.83($\pm$0.15) 4.56($\pm$0.08)
Text Prompt
SyncFusion[12] 81.6 424 - $0$ +0.529 5.11 - -
Video-Foley (Ours) 66.8 451 - 0.0136 - 0.476 3.28 - -
Text-to-Audio[6] 59.8 397 2.39($\pm$0.13) 0.0259 2.00($\pm$0.13) 0.443 2.67 2.78($\pm$0.16) 3.21($\pm$0.17)
V-B Ablation Study on the Number of RMS Bins
The number of bins for RMS discretization is a critical parameter that significantly affects both Video2RMS and RMS2Sound. To determine the optimal value, we conducted an ablation study, as presented in Figure 7. In Video2RMS, we identified a trade-off between prediction performance and computational cost, as shown on the left; more bins improve temporal synchrony but require higher complexity and more model parameters. As shown on the right, both temporal alignment performance and audio quality in RMS2Sound saturate after bins greater than 64. At 64 bins, we found no performance drop in the quantitative measures when using discretized RMS instead of continuous RMS. Therefore, we set the discretization bins to 64.
@@ V-C1 Quantitative Study
Table II compares the performance of various video-to-sound models on the GreatestHits test set. Video-Foley achieved state-of-the-art performance on all objective metrics as well as the human MOS. Notably, it showed a significant performance gap not only in temporal alignment but also in semantic material alignment compared to the audio-visual cued model (CondFoleyGen) and the onset-based model (SyncFusion). This suggests that RMS conditioning is superior for video-to-sound generation, because it conveys both timing and intensity dynamics, providing more detailed information than simple timestamps. Furthermore, this temporal feature can imply the timbre and nuance of the sound through its curve shape, complementing the semantic prompt. Diff-Foley, despite using temporal information for audio-visual joint space training, lagged in performance due to poorer temporal alignment granularity (4fps) and the frequent generation of visually irrelevant sounds, likely caused by noisy in-the-wild training sets.
@@
In every aspect, including audio quality, Video-Foley also outperforms AudioLDM [6], the frozen TTA model in RMS2Sound. This suggests that an additional RMS condition, well matched with the prompt, can help the model generate higher fidelity audio, consistent with the results in Figure 7. Video-Foley and SyncFusion, trained exclusively with audio prompts, perform better with audio prompts than text. The complexity of describing multiple sound events over 10 seconds with text versus audio may also contribute to this trend.
@@
TABLE III: Performance of Video2RMS module on different input video features. OF: Optical Flow, RGB: RGB image.
@@
V-C2 Qualitative Study
Extensive case studies were conducted to demonstrate the performance and controllability of Video-Foley. Our analysis underscores that the intensity level and energy transition in RMS are often associated with the timbre and nuance of sound, consistent with the findings of the previous study[14]. We plot the mel-spectrogram and normalized RMS of the generated audio from each model. Figure 8 shows the synergy of complex prompts with RMS. Only Video-Foley generates hit or scratch sounds at the right moment, as our model can distinguish the timing and type of each sound event from the shape of the RMS curve even for complex audio or text prompts with multiple events. Onset-based models only predict when to make a sound but cannot distinguish different timbres for each event. In contrast, ours can control both the timing and the corresponding timbre by modifying the RMS. Figure 9 illustrates the controllability and high audio-visual alignment of Video-Foley. Only ours effectively predicts and recommends the appropriate RMS level and transition curve, ensuring synchronization with the input video. This includes not only timing but also the intensity and nuance of sound events. These capabilities are due to Video2RMS’s ability to distinguish action types (e.g., hit and scratch), timing, and intensity and predict their corresponding energy transitions, and RMS2Sound’s ability to generate appropriate timbre and nuance at the corresponding timings. Additionally, RMS helps enhance temporal alignment.
V-D Ablation Study for Video2RMS
V-D1 Ablation Study on Video Features
Table III shows the objective metric scores of Video2RMS depending on the input video features. The best overall performance was achieved when both RGB and optical flow features were used. Removing either feature led to a performance drop, but excluding the optical flow resulted in a more significant decrease. This suggests that inter-frame differences captured by the optical flow are crucial for predicting temporal audio features like RMS. However, the RGB feature also enhances performance by providing semantic information, such as the presence of sound-related objects in the visual scene.
V-D2 Ablation Study on Label Smoothing
Figure 10 illustrates the performance of Video2RMS with different label smoothing window sizes $W$ in Equation 3. We found that $W=2$ offers the best balance between E-L1 and accuracy. For larger window sizes, the model qualitatively produces more jitter in the RMS curve. The prediction performance saturates after $W=10$, resulting in a spiky RMS curve and a poor overall performance. In addition, using Gaussian label smoothing ($W>0$) consistently improved performance at any window size.
@@ V-E Ablation Study for RMS2Sound
TABLE IV: Performance of RMS2Sound module. w/o RMS: pretrained AudioLDM without RMS condition, disc. RMS: discretized RMS in 64 bins, cont. RMS: continuous RMS.
@@
w/o RMS[6] 59.8 397 0.443 0.0276
disc. RMS 46.8 333 0.504 0.0050
cont. RMS 44.6 323 0.531 0.0050
Table IV summarizes the performance of RMS2Sound on audio and text prompts with ground-truth RMS conditions. The discretized RMS (64 bins) performed comparably to the original continuous RMS in terms of audio quality, semantic similarity, and temporal alignment. In contrast, the vanilla TTA model without RMS conditioning (AudioLDM) underperformed in every metric. This supports our assumption that realistic RMS conditions enhance the overall quality of the generated audio.
V-F Unlocking the Potential of RMS-ControlNet
@@
RMS-ControlNet, trained for additional temporal event guidance with RMS condition on top of the pretrained TTA model (AudioLDM), shows great potential in controllable audio generation tasks. We provide demos to showcase its high controllability, which prior TTA models were not able to achieve. Figure 11 shows how RMS can be simply and intuitively used for temporal guidance. With the same text prompt, RMS-ControlNet guides AudioLDM to generate audio that matches different input RMS conditions (A-shaped, monotonic decrease, monotonic increase, and V-shaped) that reflect varying distance from the source while maintaining audio semantics (car passing sound). Such intensity dynamics are often used in Foley sound generation, which current text-to-audio models struggle to reflect with sufficient temporal accuracy. Figure 12 shows how text prompt can adjust audio semantics along with RMS guidance. While preserving the timing of sound events, users can control various audio semantics such as the sound source and timbre. This highlights RMS-ControlNet’s ability to guarantee high controllability in RMS guidance for timing and intensity while preserving the power in TTA generation. Note that these sound sources are not part of GreatestHits; their generation leverages the knowledge embedded in the frozen TTA backbone.
VI Discussions
VI-A Future Works
Our Video-Foley model has four main limitations: two stemming from its architecture and two from the Greatest Hits dataset. Regarding the architecture, our model does not capture temporal dynamics in audio semantics, as CLAP compresses variations in sound sources and timbre into a single aggregated vector [42]. Additionally, the Video2RMS module predicts RMS solely from video input, ignoring audio or text prompts, which limits controllability — users cannot adjust sound timbre or specify sound sources. Addressing these issues requires incorporating sequential features to encode temporal changes and integrating prompts into the Video2RMS module for richer semantic guidance. Greatest Hits dataset consists of mono-sourced audio, preventing the model from handling multiple simultaneous sound sources. Furthermore, all sound sources are in the foreground, restricting spatial awareness. Expanding the dataset to include overlapping sounds and background elements is crucial for improving the model’s ability to process complex auditory scenes.
@@ VII Conclusion
VIII Acknowledgments
This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No.RS-2019-II190075, Artificial Intelligence Graduate School Program (KAIST)) and the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2023-00222383).

### v3

@@ IV-D2 Subjective Evaluation
TABLE II: Performance of the proposed Video-Foley and other video-to-sound models on Greatest Hits testset. avav: audio-video paired prompt used, +: same CLAP model for train and evaluation.
Model Audio Quality Temporal Alignment Semantic Alignment
Audio Prompt
CondFoleyGenav[20] 42.2 381 3.10($\pm$0.13) 0.148 1.93($\pm$0.13) 0.572 1.01 2.36($\pm$0.16) 2.79($\pm$0.17)
SyncFusion[12] 65.9 335 3.10($\pm$0.13) 0.150 3.10($\pm$0.19) +0.631 4.50 3.04($\pm$0.18) 3.22($\pm$0.19)
Video-Foley (Ours) 27.2 187 3.93($\pm$0.12) 0.083 4.40($\pm$0.11) 0.644 0.80 3.83($\pm$0.15) 4.56($\pm$0.08)
Text Prompt
SyncFusion[12] 81.6 424 - 0.162 - +0.529 5.11 - -
Video-Foley (Ours) 66.8 451 - 0.088 - 0.476 3.28 - -
Text-to-Audio[6] 59.8 397 2.39($\pm$0.13) 0.130 2.00($\pm$0.13) 0.443 2.67 2.78($\pm$0.16) 3.21($\pm$0.17)
Since a generated sound can be perceptually valid without exactly matching the ground truth [17], we conducted a human listening test to assess the perceptual quality of the generated audio in relation to the input video. A total of 20 participants, including audio ML researchers and audio engineers recruited via email lists and colleagues, were asked to score the audio on a five-point Likert scale based on four criteria: Material / Action / Temporal Alignment, and Audio Quality. Semantic Alignment was divided into two categories to evaluate how well the sound matches the material type and action nuance of the sound events in the video. We provided guidelines and video examples to clearly distinguish Material and Action Alignment from Temporal Alignment during evaluation. The evaluation survey consisted of 12 questions covering different material-action types. Specifically, we excluded ‘None’ from the dataset’s 18 material categories and selected six ‘{material}-scratch’ cases where the sound characteristics significantly change by scratching actions. These cases include plastic-scratch, rock-scratch, dirt-scratch, drywall-scratch, gravel-scratch, and grass-scratch. In addition, we selected six ‘{material}-hit’ cases from the remaining material categories where the sound characteristics notably change by hitting actions. These cases include carpet-hit, ceramic-hit, metal-hit, water-hit, wood-hit, and leaf-hit. To standardize the length of the sample videos and control evaluator fatigue, we trimmed each video to 4 seconds from the starting point. Each question presented the ground truth audio and the audio generated by Video-Foley, SyncFusion, Diff-Foley, CondFoleyGen, and AudioLDM in a random order. Since CondFoleyGen does not support text prompts, audio prompts were used for all models to ensure a fair comparison. The Mean Opinion Score (MOS) and its 95% confidence interval were calculated.
@@ V-A Analysis on Video2RMS
[…]
V-B Ablation Study on the Number of RMS Bins
The number of bins for RMS discretization is a critical parameter that significantly affects both Video2RMS and RMS2Sound. To determine the optimal value, we conducted an ablation study, as presented in Fig. 7. In Video2RMS, we identified a trade-off between prediction performance and computational cost, as shown on the Fig. 7(a); more bins improve temporal synchrony but require higher complexity and more model parameters. As shown in Fig. 7(b), both temporal alignment performance and audio quality in RMS2Sound saturate after bins greater than 64. At 64 bins, we found no performance drop in the quantitative measures when using discretized RMS instead of continuous RMS. Therefore, we set the discretization bins to 64.
TABLE III: Performance of the proposed Video-Foley and other single-latent video-to-sound models on Greatest Hits testset. Regarding train data, †: VGGSound[34] ($\sim$0.4k hr), $\ddagger$: VGGSound, subset of AudioSet[35] ($\sim$1.1k hr), otherwise: Greatest Hits trainset ($\sim$6 hr).
@@ V-B Ablation Study on the Number of RMS Bins
SpecVQGAN[8] 101.0 579 - 0.148 - 0.323 6.42 - -
Diff-Foley[10] 87.0 550 2.11($\pm$0.11) 0.166 1.86($\pm$0.14) 0.403 4.61 1.78($\pm$0.13) 2.38($\pm$0.17)
@@
Video-Foley (Ours) 27.2 187 3.93($\pm$0.12) 0.083 4.40($\pm$0.11) 0.644 0.80 3.83($\pm$0.15) 4.56($\pm$0.08)
@@ V-C1 Quantitative Study
Table II compares the performance of Video-Foley with other dual-latent baselines on the GreatestHits test set. Video-Foley achieved state-of-the-art performance across all objective metrics as well as the human MOS. Notably, it showed a significant performance gap not only in temporal alignment but also in semantic material alignment compared to the audio-visual cued model (CondFoleyGen) and the onset-based model (SyncFusion). This suggests that RMS conditioning is superior for video-to-sound generation, because it conveys both timing and intensity dynamics, providing more detailed information than simple timestamps. Furthermore, this temporal feature can imply the timbre and nuance of the sound through its curve shape, complementing the semantic prompt. Importantly, our model does not require timestamp annotations during training.
@@
In every aspect, including audio quality, Video-Foley also outperforms AudioLDM [6], the frozen TTA model in RMS2Sound. This suggests that an additional RMS condition, well matched with the prompt, can help the model generate higher fidelity audio, consistent with the results in Fig. 7. Video-Foley and SyncFusion, trained exclusively with audio prompts, perform better with audio prompts than text. The complexity of describing multiple sound events over 10 seconds with text versus audio may also contribute to this trend.
Table III presents the performance comparison with single-latent models to provide additional points of reference for both objective and subjective evaluations. Since these models do not use semantic prompts, they rely solely on video input for semantic alignment, putting them at a disadvantage. Diff-Foley, despite incorporating temporal information for audio-visual joint space learning, lagged in temporal performance. This may be attributed to the limited temporal alignment granularity of its video encoder (4fps) or domain mismatch between its training data and Greatest Hits, which likely led to the generation of visually irrelevant sounds common in noisy in-the-wild datasets as discussed in Section IV-C.
V-C2 Qualitative Study
Extensive case studies were conducted to demonstrate the performance and controllability of Video-Foley. Our analysis underscores that the intensity level and energy transition in RMS are often associated with the timbre and nuance of sound, consistent with the findings of the previous study[14]. We plot the mel-spectrogram and normalized RMS of the generated audio from each model. Fig. 8 shows the synergy of complex prompts with RMS. Only Video-Foley generates hit or scratch sounds at the right moment, as our model can distinguish the timing and type of each sound event from the shape of the RMS curve even for complex audio or text prompts with multiple events. Onset-based models only predict when to make a sound but cannot distinguish different timbres for each event. In contrast, ours can control both the timing and the corresponding timbre by modifying the RMS. Fig. 9 illustrates the controllability and high audio-visual alignment of Video-Foley. Only ours effectively predicts and recommends the appropriate RMS level and transition curve, ensuring synchronization with the input video. This includes not only timing but also the intensity and nuance of sound events. These capabilities are due to Video2RMS’s ability to distinguish action types (e.g., hit and scratch), timing, and intensity and predict their corresponding energy transitions, and RMS2Sound’s ability to generate appropriate timbre and nuance at the corresponding timings. Additionally, RMS helps enhance temporal alignment.
V-D Ablation Study for Video2RMS
V-D1 Ablation Study on Video Features
Table IV shows the objective metric scores of Video2RMS depending on the input video features. The best overall performance was achieved when both RGB and optical flow features were used. Removing either feature led to a performance drop, but excluding the optical flow resulted in a more significant decrease. This suggests that inter-frame differences captured by the optical flow are crucial for predicting temporal audio features like RMS. However, the RGB feature also enhances performance by providing semantic information, such as the presence of sound-related objects in the visual scene.
V-D2 Ablation Study on Label Smoothing
Fig. 10 illustrates the performance of Video2RMS with different label smoothing window sizes $W$ in Equation 3. We found that $W=2$ offers the best balance between E-L1 and accuracy. For larger window sizes, the model qualitatively produces more jitter in the RMS curve. The prediction performance saturates after $W=10$, resulting in a spiky RMS curve and a poor overall performance. In addition, using Gaussian label smoothing ($W>0$) consistently improved performance at any window size.
@@ V-D2 Ablation Study on Label Smoothing
TABLE IV: Performance of Video2RMS module on different input video features. OF: Optical Flow, RGB: RGB image.
@@ V-E Ablation Study for RMS2Sound
Table V summarizes the performance of RMS2Sound on audio and text prompts with ground-truth RMS conditions. The discretized RMS (64 bins) performed comparably to the original continuous RMS in terms of audio quality, semantic similarity, and temporal alignment. In contrast, the vanilla TTA model without RMS conditioning (AudioLDM) underperformed in every metric. This supports our assumption that realistic RMS conditions enhance the overall quality of the generated audio.
V-F Unlocking the Potential of RMS-ControlNet
RMS-ControlNet, trained for additional temporal event guidance with RMS condition on top of the pretrained TTA model (AudioLDM), shows great potential in controllable audio generation tasks. We provide demos to showcase its high controllability, which prior TTA models were not able to achieve. Fig. 11 shows how RMS can be simply and intuitively used for temporal guidance. With the same text prompt, RMS-ControlNet guides AudioLDM to generate audio that matches different input RMS conditions (A-shaped, monotonic decrease, monotonic increase, and V-shaped) that reflect varying distance from the source while maintaining audio semantics (car passing sound). Such intensity dynamics are often used in Foley sound generation, which current text-to-audio models struggle to reflect with sufficient temporal accuracy. Fig. 12 shows how text prompt can adjust audio semantics along with RMS guidance. While preserving the timing of sound events, users can control various audio semantics such as the sound source and timbre. This highlights RMS-ControlNet’s ability to guarantee high controllability in RMS guidance for timing and intensity while preserving the power in TTA generation. Note that these sound sources are not part of GreatestHits; their generation leverages the knowledge embedded in the frozen TTA backbone.
TABLE V: Performance of RMS2Sound module. w/o RMS: pretrained AudioLDM without RMS condition, disc. RMS: discretized RMS in 64 bins, cont. RMS: continuous RMS.
@@ V-F Unlocking the Potential of RMS-ControlNet
w/o RMS[6] 59.8 397 0.443 0.130
disc. RMS 46.8 333 0.504 0.037
cont. RMS 44.6 323 0.531 0.036
VI Discussions
VI-A Future Works
Our Video-Foley model has four main limitations: two stemming from its architecture and two from the Greatest Hits dataset. Regarding the architecture, our model does not capture temporal dynamics in audio semantics, as CLAP compresses variations in sound sources and timbre into a single aggregated vector [45]. Additionally, the Video2RMS module predicts RMS solely from video input, ignoring audio or text prompts, which limits controllability — users cannot adjust sound timbre temporally or specify particular sound sources. Addressing these issues requires incorporating sequential features to encode temporal changes and integrating prompts into the Video2RMS module for richer semantic guidance. Greatest Hits dataset consists of mono-sourced audio, preventing the model from handling multiple simultaneous sound sources. Furthermore, all sound sources are in the foreground, restricting spatial awareness. Expanding the dataset to include overlapping sounds and background elements is crucial for improving the model’s ability to process complex auditory scenes.
Moreover, we emphasize a broader challenge: the absence of a large-scale, high-quality video dataset that combines precise audio-visual synchrony with curated Foley sound design. While larger datasets such as VGGSound [34] ($\sim$0.4k hours) exist, their suitability for Foley sound generation is limited. As VGGSound is sourced from open-domain platforms like YouTube, it suffers from low audio-visual quality, duplicated content, and visually non-indicative sounds (e.g., off-screen or background noise) [36, 37]. This underscores the need to construct a high-quality video dataset specifically tailored for Foley applications, with carefully designed and diverse sound categories relevant to multimedia production.
@@ VI-A Future Works
Finally, we observe a lack of standardized metrics for evaluating audio-visual temporal synchrony. While E-L1, a hand-crafted distance metric, effectively captures general quality trends, small improvements in perceptual synchrony may not be reflected strictly in its value. We also considered PEAVS [41], a neural model trained to predict human opinion scores for synchrony. However, PEAVS showed no meaningful correlation with human Mean Opinion Scores (MOS), with a Spearman’s correlation of 0.49 ($p$ = 0.33 >> 0.05), compared to E-L1, which achieved a stronger correlation of -0.83 ($p$ = 0.042 << 0.05). We attribute this discrepancy to two main factors. First, PEAVS has a limited training dataset, using only 200 source videos from the AudioSet [35] evaluation split, augmented to 18.2K samples with artificial distortions. Despite the augmentation, this dataset represents only a small portion of AudioSet, which may reduce generalizability to the entire AudioSet or Greatest Hits. Second, AudioSet videos often contain poor-quality audio-visual pairs with off-screen or irrelevant sounds, potentially leading PEAVS to favor flawed outputs. These findings highlight the urgent need for a more robust and generalizable neural metric for assessing temporal synchrony in video-to-sound generation.